### v3

# Dynamics of binding three independent ligands to a single scaffold

Abstract
This note considers a system in which three ligands can independently bind to a scaffold. Such systems arise in diverse applications, including immunotherapy and synthetic biology. It is shown that there are unique steady states in each conservation class, and these are asymptotically stable. The dependency of the steady-state amount of fully bound complex, as a function of total scaffold, is analyzed as well.
Keywords: trispecific antibodies, bispecific antibodies, synthetic biology, immunology, dCAs9, CRISPR, CRN, chemical reaction networks, complex balanced, detail balanced
1 Introduction
We are interested in studying a dynamical system which describes the binding and unbinding of three ligands (let’s denote them as $\mbox{\tiny A},\mbox{\tiny B},\mbox{\tiny C}$) to a “scaffold” (denoted as Y). These ligands can independently bind to three specific sites in the scaffold, so that the possible configurations of scaffold and ligands are as shown in Figure 1.
Figure 1: Generic scaffold configurations with three possible ligands.
One motivation for studying such 3-ligand systems arises from the design of therapeutic trispecific antibodies, as shown in the diagram shown in Figure 2, left. In these applications, the “scaffold” is the trispecific antibody and the “ligands” are receptors in either cancer cells or immune cells. A different motivation comes from synthetic biology, where for example a scaffold RNA may bind to dCas9, an RNA-binding protein fused with an activation domain, and a target gene, as illustrated in Figure 2, right.
Figure 2: Left: Example of trispecific antibody, figure reproduced from [9] Right: Example of scaffold RNA and three ligands in synthetic biology CRSPR regulation, figure reproduced from [5].
We will show that, given any initial (positive) concentrations of scaffold and ligands, and no initial complexes, the dynamics of the system, assuming no further production nor elimination by degradation or dilution, is such that all concentrations converge to a uniquely defined steady state (which depends on these initial concentrations). To be more precise, the total amounts $Y_{\text{tot}}$, $A_{\text{tot}}$, $B_{\text{tot}}$, $C_{\text{tot}}$ (in any bound form) are conserved along solutions (and equal the initial concentrations of the respective unbound species). For any four such total numbers, one has what is called a stoichiometric conservation class. The main theorem states there are unique steady states in each conservation class, and these are globally asymptotically stable relative to the class. The proof relies upon showing “detailed balance” of the chemical reaction network (CRN) associated to these reactions.
In immunotherapy, it is also important to understand what is the optimal dose of antibody Y, in the sense of maximizing formation of the fully formed complex “YABC” as a function of $Y_{\text{tot}}$. The intuition is that for small amounts of $Y_{\text{tot}}$ there will be few complexes possible, but, for huge amounts of scaffold compared to ligands, ligands will rapidly bind to separate scaffolds, not allowing the opportunity for complete complexes to form. One might view this bimodality as reflecting a “competition” among the ligands for binding to the “resource” (scaffold). In the analogous “bispecific” problem in which there are just two ligands, see for example [6], this response is known to be bimodal, reaching a maximum value at exactly one value of $Y_{\text{tot}}$. We believe that a similar theorem holds for the trispecific case, and have confirmed this conjecture numerically for random values of ligand. We provide a mathematical proof only for the symmetric case: $A_{\text{tot}}=B_{\text{tot}}=C_{\text{tot}}$ and all binding and unbinding rates equal.
We remark that the independent binding assumption may not always be valid in applications. For example, in the synthetic biology example illustrated in Figure 2, promoter binding is facilitated by the previous binding of the two other components, which in effect makes the problem analogous to the bispecific antibody problem. Also, depending on the time scale of the processes, one might want to account for production and degradation/dilution of the various players (ligands and/or scaffold). These additional processes are particularly relevant in that example, and are included in the model studied in [4].
2 Mathematical preliminaries
We consider a CRN as follows. There are 11 species, denoted:
${\mathcal{S}}\;=\;\{\mbox{\tiny Y},\mbox{\tiny A},\mbox{\tiny B},\mbox{\tiny C},\mbox{\tiny YA},\mbox{\tiny YB},\mbox{\tiny YC},\mbox{\tiny YAB},\mbox{\tiny YAC},\mbox{\tiny YBC},\mbox{\tiny YABC}\}.$ (1)
We call each of A, B, C a ligand. The remaining 8 species, which all contain at least one Y, are called complexes. (We also call Y the scaffold.) We use respective lower case letters for concentrations, and arrange them into a column vector:
$x\;=\;(y,a,b,c,ya,yb,yc,yab,yac,ybc,yabc)^{\top}.$
There are 24 reactions, all with mass-action kinetics. For each ligand $\mbox{\tiny L}\in\{\mbox{\tiny A},\mbox{\tiny B},\mbox{\tiny C}\}$ and each complex $\kappa$ not already containing L, we have reactions
$\mbox{$\kappa$}+\mbox{\tiny L}\;\xrightleftharpoons[k_{\mathrm{off},\mbox{\tiny L}}]{k_{\mathrm{on},\mbox{\tiny L}}}\;\mbox{$\kappa$}\cup\{\mbox{\tiny L}\}.$
In this notation, $\mbox{$\kappa$}\cup\{\mbox{\tiny L}\}$ is shorthand for the new complex that adds the ligand L. For example, one reaction is
$\mbox{\tiny YC}+\mbox{\tiny B}\;\xrightleftharpoons[k_{\mathrm{off},\mbox{\tiny B}}]{k_{\mathrm{on},\mbox{\tiny B}}}\;\mbox{\tiny YBC}.$ (2)
A key assumption is independent binding, meaning that the rates depend only on the ligand being bound and unbound, and not on the complex. For example, another reaction is
$\mbox{\tiny YAC}+\mbox{\tiny B}\;\xrightleftharpoons[k_{\mathrm{off},\mbox{\tiny B}}]{k_{\mathrm{on},\mbox{\tiny B}}}\;\mbox{\tiny YABC}.$
(note the same $k_{\mathrm{off},\mbox{\tiny B}}$ and $k_{\mathrm{on},\mbox{\tiny B}}$). In summary, there are $12$ reversible pairs (24 total reactions), and 6 parameters $k_{\mathrm{off},\mbox{\tiny L}}$, $k_{\mathrm{on},\mbox{\tiny L}}$, $\mbox{\tiny L}\in\{\mbox{\tiny A},\mbox{\tiny B},\mbox{\tiny C}\}$.
Since all reactions are reversible, the stoichiometry matrix $\Gamma$ will be made up of pairs of columns of opposite sign. To write this matrix, Let us first introduce a matrix $\Gamma_{0}$ showing only the stoichiometry of the “forward” reactions. We order these forward reactions as follows:
$\mbox{\tiny Y}+\mbox{\tiny A}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny A}}}$ YA
$\mbox{\tiny Y}+\mbox{\tiny B}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny B}}}$ YB
$\mbox{\tiny Y}+\mbox{\tiny C}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny C}}}$ YC
$\mbox{\tiny YA}+\mbox{\tiny B}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny B}}}$ YAB
$\mbox{\tiny YB}+\mbox{\tiny A}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny A}}}$ YAB
$\mbox{\tiny YA}+\mbox{\tiny C}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny C}}}$ YAC
$\mbox{\tiny YC}+\mbox{\tiny A}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny A}}}$ YAC
$\mbox{\tiny YB}+\mbox{\tiny C}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny C}}}$ YBC
$\mbox{\tiny YC}+\mbox{\tiny B}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny B}}}$ YBC
$\mbox{\tiny YAB}+\mbox{\tiny C}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny C}}}$ YABC
$\mbox{\tiny YAC}+\mbox{\tiny B}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny B}}}$ YABC
$\mbox{\tiny YBC}+\mbox{\tiny A}$ $\xrightarrow{k_{\mathrm{on},\mbox{\tiny A}}}$ YABC
Ordering the species as in (1), the stoichiometry matrix for the forward reactions is as follows (we also show the respective vector of species concentrations):
$x=\left(\begin{array}[]{cccccccccccc}y\\ a\\ b\\ c\\ ya\\ yb\\ yc\\ yab\\ yac\\ ybc\\ yabc\end{array}\right)\quad\Gamma_{0}\;=\;\left(\begin{array}[]{cccccccccccc}-1&-1&-1&0&0&0&0&0&0&0&0&0\\ -1&0&0&0&-1&0&-1&0&0&0&0&-1\\ 0&-1&0&-1&0&0&0&0&-1&0&-1&0\\ 0&0&-1&0&0&-1&0&-1&0&-1&0&0\\ 1&0&0&-1&0&-1&0&0&0&0&0&0\\ 0&1&0&0&-1&0&0&-1&0&0&0&0\\ 0&0&1&0&0&0&-1&0&-1&0&0&0\\ 0&0&0&1&1&0&0&0&0&-1&0&0\\ 0&0&0&0&0&1&1&0&0&0&-1&0\\ 0&0&0&0&0&0&0&1&1&0&0&-1\\ 0&0&0&0&0&0&0&0&0&1&1&1\end{array}\right)$
(for example, the first column is the stoichiometry of the forward reaction $\mbox{\tiny Y}+\mbox{\tiny A}\xrightarrow{k_{\mathrm{on},\mbox{\tiny A}}}\mbox{\tiny YA}$) The complete stoichiometry matrix is $\Gamma=[\Gamma_{0},-\Gamma_{0}]$. The mass action reaction rates for the forward reactions are summarized in this vector, where a dot is used to indicate the product of concentrations:
$R_{0}(x)\;=\;\left(\begin{array}[]{cccccccccccc}k_{\mathrm{on},\mbox{\tiny A}}\,y\cdot a\\ k_{\mathrm{on},\mbox{\tiny B}}\,y\cdot b\\ k_{\mathrm{on},\mbox{\tiny C}}\,y\cdot c\\ k_{\mathrm{on},\mbox{\tiny B}}\,ya\cdot b\\ k_{\mathrm{on},\mbox{\tiny A}}\,yb\cdot a\\ k_{\mathrm{on},\mbox{\tiny C}}\,ya\cdot c\\ k_{\mathrm{on},\mbox{\tiny A}}\,yc\cdot a\\ k_{\mathrm{on},\mbox{\tiny C}}\,yb\cdot c\\ k_{\mathrm{on},\mbox{\tiny B}}\,yc\cdot b\\ k_{\mathrm{on},\mbox{\tiny C}}\,yab\cdot c\\ k_{\mathrm{on},\mbox{\tiny B}}\,yac\cdot b\\ k_{\mathrm{on},\mbox{\tiny A}}\,ybc\cdot a\end{array}\right)$
so that the full reaction vector is $R(x)=[R_{0}(x);\hat{R}_{0}(x)]$, where $\hat{R}_{0}(x)$ is obtained by replacing each $k_{\mathrm{on},\mbox{\tiny L}}$ by the respective $k_{\mathrm{off},\mbox{\tiny L}}$. The system of 24 ODE’s for the concentrations is
$\dot{x}\;=\;\Gamma R(x)\;=\;\Gamma_{0}(R_{0}(x)-\hat{R}_{0}(x))\,.$ (3)
The matrix $\Gamma_{0}$ has rank 7, and its left nullspace has a basis corresponding to the conservation laws for the scaffold and ligands:
$Y_{\text{tot}}$ $=$ $y+ya+yb+yc+yab+yac+ybc+yabc$
$A_{\text{tot}}$ $=$ $a+ya+yab+yac+yabc$
$B_{\text{tot}}$ $=$ $b+yb+yab+ybc+yabc$
$C_{\text{tot}}$ $=$ $c+yc+yac+ybc+yabc\,.$
3 Detailed balance
Consider the following matrix:
$V\;=\;\left(\begin{array}[]{cccccccccccc}-1&-1&0&0&-1\\ 1&0&-1&0&1\\ 0&1&1&0&0\\ -1&0&0&-1&-1\\ 1&0&0&0&0\\ 0&-1&0&1&0\\ 0&1&0&0&0\\ 0&0&-1&0&1\\ 0&0&1&0&0\\ 0&0&0&-1&-1\\ 0&0&0&1&0\\ 0&0&0&0&1\end{array}\right)$
and notice that this matrix has full rank 5, and that $\Gamma_{0}V=0$. Thus the columns $v_{i}$ of $V$ constitute a basis of the nullspace of $\Gamma_{0}$. Let $w_{i}:=v_{i}^{\top}$ for $i=1,2,3,4,5$. Now consider a row vector that collects all the kinetic constants, in the same order as the reactions have been listed:
$k\;:=\;(k_{\mathrm{on},\mbox{\tiny A}},k_{\mathrm{on},\mbox{\tiny B}},k_{\mathrm{on},\mbox{\tiny C}},k_{\mathrm{on},\mbox{\tiny B}},k_{\mathrm{on},\mbox{\tiny A}},k_{\mathrm{on},\mbox{\tiny C}},k_{\mathrm{on},\mbox{\tiny A}},k_{\mathrm{on},\mbox{\tiny C}},k_{\mathrm{on},\mbox{\tiny B}},k_{\mathrm{on},\mbox{\tiny C}},k_{\mathrm{on},\mbox{\tiny B}},k_{\mathrm{on},\mbox{\tiny A}})$
and use the notation $k^{w_{i}}$ to denote the pointwise powers. We have $k^{w_{i}}=1$ for each $i=1,2,3,4,5$. For example:
$k^{w_{1}}=k_{\mathrm{on},\mbox{\tiny A}}^{-1}k_{\mathrm{on},\mbox{\tiny B}}^{1}k_{\mathrm{on},\mbox{\tiny C}}^{0}k_{\mathrm{on},\mbox{\tiny B}}^{-1}k_{\mathrm{on},\mbox{\tiny A}}^{1}k_{\mathrm{on},\mbox{\tiny C}}^{0}k_{\mathrm{on},\mbox{\tiny A}}^{0}k_{\mathrm{on},\mbox{\tiny C}}^{0}k_{\mathrm{on},\mbox{\tiny B}}^{0}k_{\mathrm{on},\mbox{\tiny C}}^{0}k_{\mathrm{on},\mbox{\tiny B}}^{0}k_{\mathrm{on},\mbox{\tiny A}}^{0}\;=\;1\,.$
(A graphical interpretation of this algebraic fact is that for any two along any of cycle of reactions, if a ligand is added to the scaffold, then it must also be removed.) Finally, consider the kinetic constants corresponding to the reverse reactions:
$\tilde{k}\;:=\;(k_{\mathrm{off},\mbox{\tiny A}},k_{\mathrm{off},\mbox{\tiny B}},k_{\mathrm{off},\mbox{\tiny C}},k_{\mathrm{off},\mbox{\tiny B}},k_{\mathrm{off},\mbox{\tiny A}},k_{\mathrm{off},\mbox{\tiny C}},k_{\mathrm{off},\mbox{\tiny A}},k_{\mathrm{off},\mbox{\tiny C}},k_{\mathrm{off},\mbox{\tiny B}},k_{\mathrm{off},\mbox{\tiny C}},k_{\mathrm{off},\mbox{\tiny B}},k_{\mathrm{off},\mbox{\tiny A}})\,.$
Then also $\tilde{k}^{w_{i}}=1$ for each $i=1,2,3,4,5$. We conclude:
$k^{w_{i}}\;=\;\tilde{k}^{w_{i}}\;\;\mbox{for each }i=1,2,3,4,5\,.$ (4)
This means that the network is detailed balanced, see Remark 14.4.11 (p. 307) in Feinberg’s textbook [3]. (In the language of that reference, we have selected the forward reactions as the subset $\mathcal{R}$$^{\rightarrow}$ and the $w_{i}$ are the vectors $\alpha^{i}$.)
It follows from Theorems 14.2.1 and 14.2.3 in the same reference that each positive stoichiometric compatibility class contains precisely one equilibrium, and that this equilibrium is asymptotically stable (and in fact the linearization of the dynamics restricted to the class has its all eigenvalues real and negative).
We remark that the detailed balanced conclusion means that at each steady state each reaction is in equilibrium. For example, from the reaction in (2), we have that $k_{\mathrm{on},\mbox{\tiny B}}\,yc\cdot b=k_{\mathrm{off},\mbox{\tiny B}}\,ybc$, or, using the notation $K_{\mbox{\tiny L}}:=k_{\mathrm{on},\mbox{\tiny L}}/k_{\mathrm{off},\mbox{\tiny L}}$,
$ybc\;=\;K_{\mbox{\tiny B}}\,yc\cdot b$
and similarly for all other reactions. Inductively, it follows that, for steady state concentrations, we have
$yabc\;=\;K_{\mbox{\tiny A}}\,a\cdot ybc\;=\;K_{\mbox{\tiny A}}\,K_{\mbox{\tiny B}}\,a\cdot b\cdot yc\;=\;K_{\mbox{\tiny A}}K_{\mbox{\tiny B}}K_{\mbox{\tiny C}}\,y\cdot a\cdot b\cdot c$
and similarly for all other complexes. Thus, at steady state, the conservation laws can be rewritten, after changing to new variables $a=K_{\mbox{\tiny A}}a$, $b=K_{\mbox{\tiny B}}b$, $c=K_{\mbox{\tiny C}}c$, as follows:
$Y_{\text{tot}}$ $=$ $y\cdot\left(1+a+b+c+a\cdot b+a\cdot c+b\cdot c+a\cdot b\cdot c\right)$
$A_{\text{tot}}$ $=$ $a\cdot\left(1+y+y\cdot b+y\cdot c+y\cdot b\cdot c\right)$
$B_{\text{tot}}$ $=$ $b\cdot\left(1+y+y\cdot a+y\cdot c+y\cdot a\cdot c\right)$
$C_{\text{tot}}$ $=$ $c\cdot\left(1+y+y\cdot a+y\cdot b+y\cdot a\cdot b\right)$
and we will call these the polynomial form of the conservation laws. They represent a set of four polynomial equations in four variables, and the $(y,a,b,c)$ coordinates of the steady state constitute their unique positive solution. Conversely, the solution of these equations gives rise to the steady state of the original system once that we define $yab:=y\cdot a\cdot b$ and so forth (recall that we are assuming variables have been rescaled so that all $K_{\mbox{\tiny L}}=1$).
4 Global convergence
Recall that a subset of species $\Sigma\subseteq\mathcal{S}$ is a siphon if every reaction that produces a species in $\Sigma$ consumes at least one species in $\Sigma$:
$\forall\,y\to y^{\prime}\ \text{with}\ \operatorname{supp}(y^{\prime})\cap\Sigma\neq\varnothing:\quad\operatorname{supp}(y)\cap\Sigma\neq\varnothing.$
A siphon is critical [2] if it does not contain the support of any nonzero nonnegative conservation law (P-semiflow).
Define the moiety supports
$S_{Y}$ $=\{\mbox{\tiny Y},\mbox{\tiny YA},\mbox{\tiny YB},\mbox{\tiny YC},\mbox{\tiny YAB},\mbox{\tiny YAC},\mbox{\tiny YBC},\mbox{\tiny YABC}\},$
$S_{\mbox{\tiny A}}$ $=\{\mbox{\tiny A},\mbox{\tiny YA},\mbox{\tiny YAB},\mbox{\tiny YAC},\mbox{\tiny YABC}\},$
$S_{\mbox{\tiny B}}$ $=\{\mbox{\tiny B},\mbox{\tiny YB},\mbox{\tiny YAB},\mbox{\tiny YBC},\mbox{\tiny YABC}\},$
$S_{\mbox{\tiny C}}$ $=\{\mbox{\tiny C},\mbox{\tiny YC},\mbox{\tiny YAC},\mbox{\tiny YBC},\mbox{\tiny YABC}\}.$
Lemma 4.1 (P-semiflows and moiety siphons).
For each $\mbox{\tiny L}\in\{\mbox{\tiny Y},\mbox{\tiny A},\mbox{\tiny B},\mbox{\tiny C}\}$, the indicator vector $c^{(\mbox{\tiny L})}=\mathbf{1}_{S_{\mbox{\tiny L}}}$ satisfies $c^{(\mbox{\tiny L})\top}\Gamma=0$ (hence is a P-semiflow), and $S_{\mbox{\tiny L}}$ is a siphon.
Proof.
Each listed reaction preserves the count of each moiety: e.g., $\mbox{\tiny Y}+\mbox{\tiny A}\rightleftarrows\mbox{\tiny YA}$ preserves the number of Y’s and of A’s; similarly for all reactions. Therefore the dot product of each reaction vector with $\mathbf{1}_{S_{m}}$ is zero, giving $c^{(m)\top}\Gamma=0$. For the siphon property: in any reaction, a product contains moiety L iff some reactant contains moiety L; thus if a product lies in $S_{\mbox{\tiny L}}$, at least one reactant lies in $S_{\mbox{\tiny L}}$. Hence $S_{\mbox{\tiny L}}$ is a siphon. ∎
Lemma 4.2 (Backward-closure within a moiety).
Let $\Sigma$ be a siphon. If $\Sigma\cap S_{\mbox{\tiny L}}\neq\varnothing$ for some $\mbox{\tiny L}\in\{\mbox{\tiny Y},\mbox{\tiny A},\mbox{\tiny B},\mbox{\tiny C}\}$, then $S_{\mbox{\tiny L}}\subseteq\Sigma$.
Proof.
We give the argument for $\mbox{\tiny L}=\mbox{\tiny Y}$; the other moieties are identical.
A standard characterization says that $\Sigma$ is a siphon iff it is backward-closed under production: whenever a reaction $t\to s+\cdots$ produces some $s\in\Sigma$, at least one reactant $t$ must also lie in $\Sigma$. Therefore, starting from any $s_{0}\in\Sigma\cap S_{\mbox{\tiny Y}}$, repeatedly adding all reactants of every reaction that produces a currently included species yields the smallest siphon containing $s_{0}$.
We claim that this backward-closure from any $s_{0}\in S_{\mbox{\tiny Y}}$ generates all of $S_{\mbox{\tiny Y}}$. Indeed, enumerate the reactions that produce Y, YA, YB, YC, YAB, YAC, YBC, YABC:
  • Y is produced by $\mbox{\tiny YA}\to\mbox{\tiny Y}+\mbox{\tiny A}$, $\mbox{\tiny YB}\to\mbox{\tiny Y}+\mbox{\tiny B}$, $\mbox{\tiny YC}\to\mbox{\tiny Y}+\mbox{\tiny C}$.
  • YA is produced by $\mbox{\tiny Y}+\mbox{\tiny A}\to\mbox{\tiny YA}$, $\mbox{\tiny YAB}\to\mbox{\tiny YA}+\mbox{\tiny B}$, $\mbox{\tiny YAC}\to\mbox{\tiny YA}+\mbox{\tiny YBC}$.
  • YB is produced by $\mbox{\tiny Y}+\mbox{\tiny B}\to\mbox{\tiny YB}$, $\mbox{\tiny YAB}\to\mbox{\tiny YB}+\mbox{\tiny YA}$, $\mbox{\tiny YBC}\to\mbox{\tiny YB}+\mbox{\tiny C}$.
  • YC is produced by $\mbox{\tiny Y}+\mbox{\tiny C}\to\mbox{\tiny YC}$, $\mbox{\tiny YAC}\to\mbox{\tiny YC}+\mbox{\tiny A}$, $\mbox{\tiny YBC}\to\mbox{\tiny YC}+\mbox{\tiny B}$.
  • YAB is produced by $\mbox{\tiny YA}+\mbox{\tiny B}\to\mbox{\tiny YAB}$, $\mbox{\tiny YB}+\mbox{\tiny A}\to\mbox{\tiny YAB}$, $\mbox{\tiny YABC}\to\mbox{\tiny YAB}+\mbox{\tiny YC}$.
  • YAC is produced by $\mbox{\tiny YA}+\mbox{\tiny C}\to\mbox{\tiny YAC}$, $\mbox{\tiny YC}+\mbox{\tiny A}\to\mbox{\tiny YAC}$, $\mbox{\tiny YABC}\to\mbox{\tiny YAC}+\mbox{\tiny B}$.
  • YBC is produced by $\mbox{\tiny YB}+\mbox{\tiny C}\to\mbox{\tiny YBC}$, $\mbox{\tiny YC}+\mbox{\tiny B}\to\mbox{\tiny YBC}$, $\mbox{\tiny YABC}\to\mbox{\tiny YBC}+\mbox{\tiny A}$.
  • YABC is produced by $\mbox{\tiny YAB}+\mbox{\tiny C}\to\mbox{\tiny YABC}$, $\mbox{\tiny YAC}+\mbox{\tiny B}\to\mbox{\tiny YABC}$, $\mbox{\tiny YBC}+\mbox{\tiny A}\to\mbox{\tiny YABC}$.
From any starting $s_{0}\in S_{\mbox{\tiny Y}}$, repeatedly including all reactants of the productions above forces inclusion of every Y-carrying reactant listed, hence eventually of every Y-carrying species. Concretely, if $s_{0}\in\{\mbox{\tiny YA},\mbox{\tiny YB},\mbox{\tiny YC}\}$, the productions of Y force inclusion of the other two among $\{\mbox{\tiny YA},\mbox{\tiny YB},\mbox{\tiny YC}\}$; then the productions of $\mbox{\tiny YA},\mbox{\tiny YB},\mbox{\tiny YC}$ force inclusion of $\mbox{\tiny YAB},\mbox{\tiny YAC},\mbox{\tiny YBC}$; then productions of $\mbox{\tiny YAB},\mbox{\tiny YAC},\mbox{\tiny YBC}$ force inclusion of YABC; conversely, productions of $\mbox{\tiny YAB},\mbox{\tiny YAC},\mbox{\tiny YBC}$ from YABC do not remove anything. If $s_{0}$ is already among $\{\mbox{\tiny YAB},\mbox{\tiny YAC},\mbox{\tiny YBC},\mbox{\tiny YABC}\}$, the same backward-closure still adds the remaining Y-carriers via the bullets above. Thus the minimal backward-closed set containing $s_{0}$ is $S_{\mbox{\tiny Y}}$, and since a siphon is backward-closed, $S_{\mbox{\tiny Y}}\subseteq\Sigma$. ∎
Corollary 1 (Structure of siphons).
Every siphon $\Sigma$ is a union of moiety supports:
$\exists\,L\subseteq\{\mbox{\tiny Y},\mbox{\tiny A},\mbox{\tiny B},\mbox{\tiny C}\}\quad\text{such that}\quad\Sigma=\bigcup_{\mbox{\tiny L}\in L}S_{\mbox{\tiny L}}.$
In particular, the minimal nonempty siphons are exactly $S_{\mbox{\tiny Y}},S_{\mbox{\tiny A}},S_{\mbox{\tiny B}},S_{\mbox{\tiny C}}$.
Proof.
Let $L=\{\mbox{\tiny L}:\Sigma\cap S_{\mbox{\tiny L}}\neq\varnothing\}$. By Lemma 4.2, $S_{\mbox{\tiny L}}\subseteq\Sigma$ for each $\mbox{\tiny L}\in L$, hence $\bigcup_{\mbox{\tiny L}\in L}S_{\mbox{\tiny L}}\subseteq\Sigma$. The reverse inclusion is immediate from the definition of $L$. Minimality follows: if $S_{\mbox{\tiny L}}$ properly contained a smaller nonempty siphon, that smaller siphon would intersect $S_{\mbox{\tiny L}}$ and therefore equal $S_{\mbox{\tiny L}}$ by Lemma 4.2. ∎
Corollary 2 (No critical siphons).
The CRN has no nonempty critical siphons.
Proof.
By Lemma 4.1, for each $\mbox{\tiny L}\in\{\mbox{\tiny Y},\mbox{\tiny A},\mbox{\tiny B},\mbox{\tiny C}\}$ the indicator $c^{(\mbox{\tiny L})}$ is a P-semiflow with support $S_{\mbox{\tiny L}}$. By Corollary 1, every minimal nonempty siphon is one of the $S_{\mbox{\tiny L}}$, hence contains the support of a P-semiflow. Therefore no siphon is critical. ∎
Theorem 4.3 (Global stability).
The unique positive steady state in each stoichiometry class is globally asymptotically stable with respect to the class.
Proof.
To prove Theorem 4.3, we remark that if a CRN has no critical siphons (every siphon contains the support of a P-semiflow), then all trajectories with strictly positive initial condition are persistent, meaning that they do not approach the boundary [2]. It is also known that or mass-action systems admitting the standard Horn–Jackson Lyapunov function, and in particular for complex-balanced systems, trajectories either converge to the interior equilibrium or to the boundary (see e.g. [8, 1]). As persistence rules out the boundary alternative, we conclude global asymptotic stability of the (unique) complex-balanced equilibrium in each positive stoichiometric compatibility class. ∎
Remark $4.$
An alternative proof would be based on showing that there are no steady states in the boundary of any positive stoichiometric class, and that complex balanced systems with no boundary equilibria have global asymptotic stability in each class [7, 8].
5 Steady state of YABC as a function of $Y_{\text{tot}}$
As explained in the introduction, one is often interested in the steady state concentration yabcyabc of the full complex as a function of the “dose” $Y_{\text{tot}}$. Assuming that we have fixed positive values for the total concentrations $(A_{\text{tot}},B_{\text{tot}},C_{\text{tot}})$ of ligands, and that we write $r:=Y_{\text{tot}}$, this function is:
$f(r)=y(r)\cdot a(r)\cdot b(r)\cdot c(r)$
where $(y(r),a(r),b(r),c(r)$ is the unique positive solution of the polynomial form of the conservation laws. It is not difficult, using the implicit function theorem, to show that $f$ is differentiable, and in fact $\mathcal{C}$$^{\infty}$.
Note that, at steady states, the conservation laws imply that $y(r)<r=Y_{\text{tot}}$ and
$a(r)\;<\;\frac{A_{\text{tot}}}{1+y(r)},\;\;b(r)\;<\;\frac{B_{\text{tot}}}{1+y(r)},\;\;c(r)\;<\;\frac{C_{\text{tot}}}{1+y(r)},$
so
$f(r)\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{y(r)}{(1+y(r))^{3}}\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1+y(r)}{(1+y(r))^{3}}\;=\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1}{(1+y(r))^{3}}$ (5)
for all $r>0$. So, as $\lim_{r\rightarrow 0}y(r)=0$, and since $a,b,c$ are bounded it also follows that $\lim_{r\rightarrow 0}f(r)=0$. Furthermore,
$y(r)\;=\;\frac{Y_{\text{tot}}}{1+a(r)+b(r)+c(r)+a(r)\cdot b(r)+a(r)\cdot c(r)+b(r)\cdot c(r)+a(r)\cdot b(r)\cdot c(r)}\;>\;\frac{Y_{\text{tot}}}{M}$
where $M=1+A_{\text{tot}}+\ldots$ is a polynomial on $A_{\text{tot}},B_{\text{tot}},C_{\text{tot}}$, because each ligand is upper bounded by its total. Thus $\lim_{r\rightarrow\infty}y(r)=\infty$, and from here and (5) we conclude that $\lim_{r\rightarrow\infty}f(r)=0$ as well. See Figure 3 for an illustration.
Figure 3: Example of a typical function $f(r)$. The parameters here are: $(A_{\text{tot}},B_{\text{tot}},C_{\text{tot}})=(1,5,10)$ and (without loss of generality, after rescaling) $K_{\mbox{\tiny A}}=K_{\mbox{\tiny B}}=K_{\mbox{\tiny C}}=1$.
Numerical simulations suggest the following, for all parameters:
Conjecture 1.
The dose function $f$ is biphasic: there is an $r_{0}>0$ such that $f$ is strictly increasing for $r\in(0,r_{0})$ and strictly decreasing for $r\in(r_{0},\infty)$.
5.1 The symmetric case
Next, we prove this property only for the special symmetric case in which $A_{\text{tot}}=B_{\text{tot}}=C_{\text{tot}}$. Clearly, by symmetry, the unique steady state has $a=b=c$. Thus, the goal is to show that, for the unique solution of
$y(1+a)^{3}$ $=$ $r$
$a(1+y(1+a)^{2})$ $=$ $p$
where $p>0$ is fixed, and $r>0$ varies (we dropped the dots since there is no ambiguity now) it follows that for some $r_{0}>0$, $f(r)=ya^{3}$ is strictly increasing for $r\in(0,r_{0})$ and strictly decreasing for $r\in(r_{0},\infty)$. Since the first equation allows us to eliminate $y=r/(1+a)^{3}$, the following result is what we need.
Proposition 1.
Fix a constant $p>0$. For each $r>0$, let $a=\alpha(r)>0$ be the unique solution of the implicit equation
$a\!\left(1+\frac{r}{1+a}\right)=p.$ (6)
Define
$f(r)=\frac{\alpha(r)^{3}\,r}{\bigl{(}1+a(r)\bigr{)}^{3}}\quad\mbox{for}\;\;r>0.$ (7)
Then $f$ possesses exactly one local (and hence global) maximum on the interval $(0,\infty)$.
Proof.
We first remark that we know that a unique positive solution of (6) exists, because this is a special case of the general theory. However, we can also see this directly because, ignoring the positive denominator $1+a$, the equation can be written as
$a^{2}+(r-p+1)a-p\;=\;0$
which is a parabola that is negative at $a=0$ and =pr>0=pr>0 at $a=p$. From the quadratic formula, it is clear that $\alpha$ is differentiable. Now define:
$\rho(a):=\frac{(p-a)\,(a+1)}{a}\quad\mbox{for}\;\;0<a<p$
and notice that $r=\rho(a)$ if and only if $a=\alpha(r)$, i.e. $\rho$ is the inverse of $\alpha$, $\alpha(\rho(a))=a$ for all $0<a<p$. The derivative
$\rho^{\prime}(a)\;=\;-\bigl{(}1+p/a^{2}\bigr{)}<0$
shows that the function $\rho$ is strictly decreasing and smooth on $(0,p)$. Define the function:
$g(a):=f\!\bigl{(}\rho(a)\bigr{)},\qquad a\in(0,p).$
From the chain rule, $g^{\prime}(a)=f^{\prime}(\rho(a))\rho^{\prime}(a)$, and thus
$f^{\prime}(r)\;=\;\frac{g^{\prime}(\alpha(r))}{\rho^{\prime}(\alpha(r))},\qquad r\in(0,\infty).$
We had seen that the denominator is always negative. Thus, if we prove that for some $a^{*}\in(0,p)$, $g^{\prime}(a)>0$ on $(0,a^{*})$ and $g^{\prime}(a)<0$ on $(a^{*},\infty)$, it follows that for $r^{*}=\rho(a^{*})\in(0,p)$, $f^{\prime}(r)>0$ on $(0,r^{*})$ and $f^{\prime}(r)<0$ on $(r^{*},\infty)$. (Note that, since $\alpha$ decreases, $(0,a^{*})$ gets mapped onto $(r^{*},\infty)$ and $(a^{*},p)$ gets mapped onto $(0,r^{*})$.) Using that $\alpha(\rho(a))=a$, we have that
$g(a)\;=\;a^{3}\,\frac{\rho(a)}{(1+a)^{3}}=\frac{a^{3}}{(1+a)^{3}}\,\frac{(p-a)(a+1)}{a}=\frac{a^{2}\,(p-a)}{(1+a)^{2}}$
and thus
$g^{\prime}(a)\;=\;-\frac{a}{(a+1)^{3}}P(a)\,,\quad P(a)\;:=\;a^{2}+3a-2p\,.$
Because of the negative sign and the fact that the first fraction is positive, all we need to show is that, for some $a^{*}\in(0,p)$, $P(a)<0$ for $a\in(0,a^{*})$ and $P(a)>0$ for $a\in(a^{*},p)$. But this is clear because the parabola $P(a)$ is negative at $a=0$ and positive ($p^{2}+p$) at $a=p$. ∎
For completeness, let us also analyze the bispecific case.
5.2 The bispecific case
We show here that, for the bispecific problem, in which the species are $y,a,b$, the steady-state amount yabyab of fully formed complexes has a unique local (and hence global) maximum as a function of $Y_{\text{tot}}$. As with the trispecific case, this system is also detailed balanced, so that $yab=y\cdot a\cdot b$. Thus, we want to show, writing $a_{1}:=a$, $a_{2}:=b$, $r:=Y_{\text{tot}}$, $p_{1}:=A_{\text{tot}}$, $p_{2}:=B_{\text{tot}}$, and dropping dots for notational simplicity:
Theorem 5.1.
Consider the following set of algebraic equations on $(y,a_{1},a_{2})$:
$y(1+a_{1})(1+a_{2})$ $=$ $r$ (8)
$a_{1}(1+y(1+a_{2}))$ $=$ $p_{1}$ (9)
$a_{2}(1+y(1+a_{1}))$ $=$ $p_{2}\,.$ (10)
For any positive $(r,p_{1},p_{2})$, there is a unique positive solution, which is a smooth function of $(r,p_{1},p_{2})$. Now fix any positive $(p_{1},p_{2})$, write the solution $(y(r),a_{1}(r),a_{2}(r))$ as a function of $r$, and consider the function
$f(r)\,:=\;y(r)\,a_{1}(r)\,a_{2}(r)$
for $r>0$. Then $f$ has a unique local (and hence global) maximum.
Existence and uniqueness follow by the same argument as in the trispecific case (or could be proved directly). Smoothness follows by the implicit function theorem, as the Jacobian of the left-hand side has determinant $(1+a_{1}+y+ya_{2})(1+a_{2}+y+ya_{1})$, which is nonzero (in fact, positive) for all $(y,a_{1},a_{2})$. As in the trispecific case, $\lim_{r\rightarrow 0}f(r)=\lim_{r\rightarrow\infty}f(r)=0$. Thus we only need to show that the derivative $f^{\prime}(r)$ can vanish at most once.
Now consider the equations (9)-(10), with the $p_{i}$’s fixed, as a set of equations on $(a_{1},a_{2})$ with $y$ thought of as a positive parameter. Eliminating $a_{1}=\frac{p_{1}}{1+y+ya_{2}}$ from (9) and substituting into (10) we obtain $n/d=0$, where $d=1+y+ya_{2}$ and we must solve
$n\;=\;\mu a_{2}^{2}+\beta a_{2}-\kappa\;=\;0$
where
$\mu=y^{2}+y>0,\;\beta=1+2y+p_{1}y-p_{2}y+y^{2},\;\kappa=p_{2}+p_{2}y>0\,.$
This has a unique positive solution $a_{2}=\alpha_{2}(y)=\frac{-\beta+\sqrt{\beta^{2}+4\mu\kappa}}{2\mu}$, which is smooth on $y$, and thus also $a_{1}=\alpha_{1}(y)=\frac{p_{1}}{1+y+y\alpha_{2}(y)}$ is uniquely defined and smooth. We show below that $\alpha_{i}^{\prime}(y)<0$ for all $y$, $i=1,2$. Assuming this for now, define the function $\rho(y):=y(1+\alpha_{1}(y))(1+\alpha_{2}(y))$ and observe that, from equation (9) we know that
$y(1+\alpha_{2}(y))\;=\;\frac{p_{1}}{\alpha_{1}(y)}-1$
so
$\rho(y)\;=\;(1+\alpha_{1}(y))\left(\frac{p_{1}}{\alpha_{1}(y)}-1\right)\;=\;p_{1}-1-\alpha_{1}(y)+\frac{p_{1}}{\alpha_{1}(y)}$
and by the chain rule
$\rho^{\prime}(y)\;=\;-\left(1+\frac{p_{1}}{(\alpha_{1}(y))^{2}}\right)\,\alpha_{1}^{\prime}(y)\;>\;0$
(strictly increasing $\rho$). Therefore $\rho(y)=r$ if and only if $y=\eta(r)$, where $\eta$ is the inverse function of $\rho$, and $\eta$ is also strictly increasing. Now consider the function
$\varphi(y)\,:=\;y\,\alpha_{1}(y)\,\alpha_{2}(y)$
and note that $f(r)=\varphi(\eta(r))$. Therefore $f^{\prime}(r)=\varphi^{\prime}(\eta(r))\eta^{\prime}(r)$. As $\eta^{\prime}(r)\not=0$ for all $r$ (and it is one to one), we are reduced to showing that $\varphi^{\prime}$ has only one zero. This is proved next.
Lemma 5.2.
Consider the following set of algebraic equations on $(a_{1},a_{2})$:
$a_{1}(1+y(1+a_{2}))$ $=$ $p_{1}$
$a_{2}(1+y(1+a_{1}))$ $=$ $p_{2}\,.$
For any positive $(y,p_{1},p_{2})$, there is a unique positive solution, which is a smooth function of $(y,p_{1},p_{2})$. Now fix any positive $(p_{1},p_{2})$, write the solution $(a_{1},a_{2})=(\alpha_{1}(y),\alpha_{2}(y))$ as a function of $y$, and consider
$\varphi(y)\,:=\;y\,\alpha_{1}(y)\,\alpha_{2}(y)$
as a function of $y>0$. Then $\varphi^{\prime}(y)=0$ has at most one solution, and $\alpha_{i}$ is strictly decreasing, $i=1,2$.
Proof.
Introduce
$h(t)\,:=\;\log\varphi(e^{t}),\;t\in{\mathbb{R}}\,.$
We will show that $h$ is strictly concave, $h^{\prime\prime}(t)<0$ for all $t\in{\mathbb{R}}$. It then follows that $h^{\prime}(t)=0$ has at most one solution, and therefore also $\varphi^{\prime}(y)=0$ has at most one solution:
$h^{\prime}(t)\;=\;\frac{e^{t}}{\varphi(e^{t})}\,\varphi^{\prime}(e^{t})\,$,
so, since the fraction is always positive, the numbers of zeros of $h^{\prime}$ and $\varphi^{\prime}$ are the same. We use implicit differentiation to find the sign of $h^{\prime\prime}$. Consider the change of variables $t=\log y$ and the functions
$\eta_{i}(t)\,:=\;\log\alpha_{i}(e^{t})\,\;i=1,2\,.$
In these variables,
$h(t)\;=\;t+z_{1}(t)+z_{2}(t)\,.$
Introduce the function $G:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}^{2}$:
$G(t,z_{1},z_{2})\,:=\;\left(\begin{array}[]{cccccccccccc}z_{1}+\log\left(1+e^{t}\left(1+e^{z_{2}}\right)\right)\cr z_{2}+\log\left(1+e^{t}\left(1+e^{z_{1}}\right)\right)\end{array}\right)$
so that the equations are summarized as
$G(t,\eta_{1}(t),\eta_{2}(t))=(\log p_{1},\log p_{2})^{\top}\,.$
Let $G_{z}$ denote the Jacobian of $G$ with respect to $(z_{1},z_{2})$ and $G_{t}$ the derivative with respect to $t$. These and a $2\times 2$ matrix and a 2-vector respectively, and implicit differentiation gives (writing $\eta$ for the vector with coordinates $\eta_{1},\eta_{2})$:
$G_{t}(t,\eta(t))+G_{z}(t,\eta(t))\eta^{\prime}(t)\;\equiv\;0$
Now, after substituting back $t=y$ and $z_{i}=\log a_{i}$, we have that
$\det G_{z}\;=\;\frac{(1+y)(1+y+ya_{1}+ya_{2})}{(1+y+ya_{1})(1+y+ya_{2})}$
is always nonzero, so
$\eta^{\prime}(t)\;=\;-\left(G_{z}(t,\eta(t))\right)^{-1}G_{t}(t,\eta(t))\,.$
Calculating explicitly, we obtain $\eta_{1}^{\prime}(t)=n_{1}(t)/d_{1}(t)$, where
$n_{1}(t)\;=\;-e^{t}(1+e^{t}+e^{z_{2}}+e^{t}e^{z_{1}}),\quad d_{1}(t)\;=\;(1+e^{t})(1+e^{t}e^{z_{1}}+e^{t}e^{z_{2}})$
and similarly for $\eta_{2}^{\prime}$, so both functions $\eta_{i}$ are strictly decreasing. Since $\log$ and exponential are strictly increasing and $\eta_{i}(t)=\log\alpha_{i}(e^{t})$, it follows that the functions $\alpha_{i}$ are strictly decreasing.
Taking derivatives one again, and now omitting arguments for legibility:
$G_{tt}+2G_{tz}\eta^{\prime}+(\eta^{\prime})^{\top}H\eta+G_{z}\eta^{\prime\prime}\;\equiv\;0$
where $G_{tz}$ a matrix of mixed partial derivatives, and $H$ is the $2\times 2$ matrix:
$H\,:=\;\left(\begin{array}[]{cccccccccccc}(G_{1})_{zz}&0\cr 0&(G_{2})_{zz}\end{array}\right)$
where $(G_{i})_{zz}$ is the Hessian of the $i$th component of $G$. Thus
$\eta^{\prime\prime}(t)\;=\;-\left(G_{z}(t,\eta(t))\right)^{-1}\left(G_{tt}+2G_{tz}\eta^{\prime}+(\eta^{\prime})^{\top}H\eta\right)\,.$
Observe that $h^{\prime\prime}(t)=\eta_{1}^{\prime\prime}(t)+\eta_{2}^{\prime\prime}(t)$. An explicit computation gives, after substituting back $t=y$ and $z_{i}=\log a_{i}$, that $h^{\prime\prime}=n/d$, where
$n\;=\;-y(a_{1}^{2}y^{3}+a_{1}^{2}y+4ya_{1}a_{2}^{2}+ya_{1}^{3}+3ya_{1}^{2}+3ya_{1}+a_{1}+a_{2}^{2}y^{3}+a_{2}^{2}y+ya_{2}^{3}+3ya_{2}^{2}+3ya_{2}+a_{2}+2y^{2}+4y+2)$
and
$d\;=\;(1+y+1)(1+y+ya_{1}+ya_{2})^{3}\,.$
As all terms are positive, it follows that $h^{\prime\prime}<0$ for all $t$, as desired. ∎
Remark 5.3.
Unfortunately, the same proof fails in the trispecific case. Indeed, for suitable choices of $y,a_{i}$, one obtains that $h^{\prime\prime}(t)>0$ (and $h^{\prime\prime}(t)<0$ for other values of $y$). For example, with $(a_{1},a_{2},a_{3})=(10,0.1,0.1)$ one obtains $h^{\prime\prime}$ negative, positive, and again negative, at y=0.5, 1, 4y=0.5,\,1,\,4 respectively. Thus there are at least two points where $h^{\prime\prime}=0$ (roughly at $y=0.6995$ and $3.5068$).
Acknowledgement
This work was partially supported by NSF grant DMS-2052455 and AFOSR grants FA-9550-21-1-0289 and FA9550-22-1-0316.

### v4

@@
 and similarly for all other complexes.
-Thus, at steady state, the conservation laws can be rewritten, after changing to new variables $a=K_{\mbox{\tiny A}}a$, $b=K_{\mbox{\tiny B}}b$, $c=K_{\mbox{\tiny C}}c$, as follows:
-$Y_{\text{tot}}$ $=$ $y\cdot\left(1+a+b+c+a\cdot b+a\cdot c+b\cdot c+a\cdot b\cdot c\right)$
-$A_{\text{tot}}$ $=$ $a\cdot\left(1+y+y\cdot b+y\cdot c+y\cdot b\cdot c\right)$
-$B_{\text{tot}}$ $=$ $b\cdot\left(1+y+y\cdot a+y\cdot c+y\cdot a\cdot c\right)$
-$C_{\text{tot}}$ $=$ $c\cdot\left(1+y+y\cdot a+y\cdot b+y\cdot a\cdot b\right)$
-and we will call these the polynomial form of the conservation laws.
-They represent a set of four polynomial equations in four variables, and the $(y,a,b,c)$ coordinates of the steady state constitute their unique positive solution.
-Conversely, the solution of these equations gives rise to the steady state of the original system once that we define $yab:=y\cdot a\cdot b$ and so forth (recall that we are assuming variables have been rescaled so that all $K_{\mbox{\tiny L}}=1$).
 4 Global convergence
@@
 5 Steady state of YABC as a function of $Y_{\text{tot}}$
+We will assume for simplicity that $K_{\mbox{\tiny A}}=K_{\mbox{\tiny B}}=K_{\mbox{\tiny C}}=1$ and leave the general case for a revised version.
+Thus, at steady state, the conservation laws can be rewritten as follows:
+$Y_{\text{tot}}$ $=$ $y\cdot\left(1+a+b+c+a\cdot b+a\cdot c+b\cdot c+a\cdot b\cdot c\right)$
+$A_{\text{tot}}$ $=$ $a\cdot\left(1+y+y\cdot b+y\cdot c+y\cdot b\cdot c\right)$
+$B_{\text{tot}}$ $=$ $b\cdot\left(1+y+y\cdot a+y\cdot c+y\cdot a\cdot c\right)$
+$C_{\text{tot}}$ $=$ $c\cdot\left(1+y+y\cdot a+y\cdot b+y\cdot a\cdot b\right)$
+and we will call these the polynomial form of the conservation laws.
+They represent a set of four polynomial equations in four variables, and the $(y,a,b,c)$ coordinates of the steady state constitute their unique positive solution.
+Conversely, the solution of these equations gives rise to the steady state of the original system once that we define $yab:=y\cdot a\cdot b$ and so forth.
 As explained in the introduction, one is often interested in the steady state concentration yabcyabc of the full complex as a function of the “dose” $Y_{\text{tot}}$.
@@
 so
-$f(r)\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{y(r)}{(1+y(r))^{3}}\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1+y(r)}{(1+y(r))^{3}}\;=\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1}{(1+y(r))^{3}}$ (5)
+$f(r)\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{y(r)}{(1+y(r))^{3}}\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1+y(r)}{(1+y(r))^{3}}\;=\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1}{(1+y(r))^{2}}$ (5)
 for all $r>0$. So, as $\lim_{r\rightarrow 0}y(r)=0$, and since $a,b,c$ are bounded it also follows that $\lim_{r\rightarrow 0}f(r)=0$. Furthermore,
@@
 See Figure 3 for an illustration.
-Figure 3: Example of a typical function $f(r)$.
-The parameters here are: $(A_{\text{tot}},B_{\text{tot}},C_{\text{tot}})=(1,5,10)$ and (without loss of generality, after rescaling) $K_{\mbox{\tiny A}}=K_{\mbox{\tiny B}}=K_{\mbox{\tiny C}}=1$.
+Figure 3: Example of a typical function $f(r)$. The parameters here are: $(A_{\text{tot}},B_{\text{tot}},C_{\text{tot}})=(1,5,10)$ and $K_{\mbox{\tiny A}}=K_{\mbox{\tiny B}}=K_{\mbox{\tiny C}}=1$.
 Numerical simulations suggest the following, for all parameters:
@@
 Existence and uniqueness follow by the same argument as in the trispecific case (or could be proved directly).
-Smoothness follows by the implicit function theorem, as the Jacobian of the left-hand side has determinant $(1+a_{1}+y+ya_{2})(1+a_{2}+y+ya_{1})$, which is nonzero (in fact, positive) for all $(y,a_{1},a_{2})$.
+(See also [11, 4] for other proofs, including non-detailed balanced cases.) Smoothness follows by the implicit function theorem, as the Jacobian of the left-hand side has determinant $(1+a_{1}+y+ya_{2})(1+a_{2}+y+ya_{1})$, which is nonzero (in fact, positive) for all $(y,a_{1},a_{2})$.
 As in the trispecific case, $\lim_{r\rightarrow 0}f(r)=\lim_{r\rightarrow\infty}f(r)=0$.
@@
 Thus there are at least two points where $h^{\prime\prime}=0$ (roughly at $y=0.6995$ and $3.5068$).
+Remark 5.4.
+One can show that the following functions of $(r,p_{1},p_{2})$ solve the equations (8)-(10). For convenience, introduce
+$P_{1}=\sqrt{(r-p_{1}-1)^{2}+4r}$
+and
+$P_{2}=\sqrt{(r-p_{2}-1)^{2}+4r}\,.$
+Then, at steady state:
+$y$ $=$ $\frac{(p_{1}-r+1-P_{1})\,(p_{2}-r+1-P_{2})}{4r}$
+$a_{1}$ $=$ $\frac{p_{1}-r-1+P_{1}}{2}$
+$a_{2}$ $=$ $\frac{p_{2}-r-1+P_{2}}{2}\,.$
+These formulas can be verified by substitution. Furthermore, one can algebraically show, after simplification, that
+$y\cdot a_{1}\cdot a_{2}\;=\;\frac{1}{4r}\,(r+p_{1}+1-P_{1})(r+p_{2}+1-P_{2}).$
+This last formula can be easily shown to be the same as the one obtained in [4], when $K_{\mbox{\tiny A}}=K_{\mbox{\tiny B}}=1$.
+(The formula for $y$ would match, after simplifications, eqn.
+(S57) (for “[B]”) in that reference but only if there is an $r$ instead of $r^{2}$ in the denominator.)
 Acknowledgement